OpenStudy (anonymous):

solve 6x+4/4x-3=3x+2/2x+5 A;-2/3 B;-13/5 C;{}

OpenStudy (anonymous):

@keith15

OpenStudy (anonymous):

@Yahoo! I NEED HELP PLEASE

OpenStudy (anonymous):

6x+4/4x-3=3x+2/2x+5 Collect the like terms and solve

OpenStudy (anonymous):

What is A, B and C here??

OpenStudy (anonymous):

THE ANSWER

OpenStudy (anonymous):

@YAHOO NO SOLUCTION

OpenStudy (anonymous):

6x+4/4x-3=3x+2/2x+5 6x - 3x +4/4x - 2/2x = 8 3x + (4-4)/4x = 8 3x + 0 = 8 SO x ???

OpenStudy (anonymous):

NO SOLUTION

OpenStudy (anonymous):

@gerryliyana NO SOLUTION

OpenStudy (anonymous):

rewrite use equation tab please

OpenStudy (anonymous):

A;-2/3

OpenStudy (anonymous):

A;-2/3

OpenStudy (anonymous):

@ajprincess HELP ME SAME OF ROBERT 5

OpenStudy (anonymous):

Are they options given??

OpenStudy (anonymous):

@ajprincess HELP PLEASE;)

OpenStudy (anonymous):

;)

OpenStudy (ajprincess):